There are 10 white and 5 black balls in the box. Four balls one after the other (returning ball to
the box) were taken from the box. Let X is the number of black ball selected.
a) Find a probability density function f(x) and a cumulative distributional function F(x);
b) What is the mean and the variance?
c) What is the probability that exactly 2 balls will be white?
1
Expert's answer
2020-04-08T09:53:39-0400

"p=\frac{1}{3}, \, q=1-p," and "n=4"

(a). "f(x)=\binom{n}{x}×p^x×q^{n-x}, x=0,1,2,3,4,"

"F(x,i)=\sum_{i}^{x}\binom{n}{i}p^iq^{n-i}"

b. "E(x)=np=\frac{4}{3}=1\frac{1}{3}"

"Var(x)=npq=\frac{4}{3}×\frac{2}{3}=\frac{8}{9}"

c. from a, we have "f(2)=\binom{4}{2}×(\frac{1}{3})^2×(\frac{2}{3})^{4-2}"

"=6×\frac{1}{9}×\frac{4}{9}=\frac{24}{81}"